A poll is given, showing 35% are in favor of a new building project.

If 8 people are chosen at random, what is the probability that exactly 1 of them favor the new building project?
Answer #1

Here, n = 8, p = 0.35, (1 - p) = 0.65 and x = 1
As per binomial distribution formula P(X = x) = nCx * p^x * (1 - p)^(n - x)

We need to calculate P(X = 1)
P(X = 1) = 8C1 * 0.35^1 * 0.65^7
P(X = 1) = 0.1373